# PoseBench3D: A Cross-Dataset Analysis Framework for 3D Human Pose Estimation

### v1

@@ 4.3 Impact of Z-Score Standardization
Z-score standardization is a common operation to normalize the data by zero-centering with the mean and scaling by the standard deviation. In the context of pose estimation, where datasets may originate from diverse distributions, such normalization indeed improves cross-dataset performance when the mean and standard deviation are known—though they are typically unavailable during inference time in real-world scenarios. Table 4 shows the models [22] and [42] trained with H36M and normalized with the mean and standard deviation of the same dataset. For the other datasets, we used their respective mean and standard deviation to normalize them. We also include the unoptimized ($\diamond$) and PoseAug-optimized ($\dagger$) variants from table 2 for comparison, and also note the percentage improvement relative to the unoptimized variant in both cases. From all the cases, note that z-score normalization performed with the test set’s mean and standard deviation improves cross-dataset generalization significantly for both models (e.g., SemGCN sees an improvement of more than 50% in 3DPW). PoseAug’s optimization adds novel poses to the training and secures a place right in the middle of the unoptimized variant and the z-score standardized variant. This shows that models overfit on the distribution they are trained on—which can often be countered by increasing variation.
Table 4: MPJPE comparison of Martinez and SEM-GCN models across datasets. Percentage improvements relative to the baseline ($\diamond$) are in parentheses.
@@
Unoptimised $\diamond$ 41.42 205.62 226.20 110.01 145.81
Optimised $\dagger$ 39.11 ($\downarrow$ 5.6%) 169.79 ($\downarrow$ 17.4%) 134.12 ($\downarrow$ 40.7%) 98.99 ($\downarrow$ 10.0%) 110.50 ($\downarrow$ 24.2%)
Z-score 52.37 ($\uparrow$ 26.4%) 104.39 ($\downarrow$ 49.3%) 141.10 ($\downarrow$ 37.6%) 81.64 ($\downarrow$ 25.8%) 94.88 ($\downarrow$ 34.9%)
@@
Unoptimised $\diamond$ 47.03 262.34 315.31 118.30 185.75
Optimised $\dagger$ 41.90 ($\downarrow$ 10.9%) 241.21 ($\downarrow$ 8.1%) 239.07 ($\downarrow$ 24.2%) 107.26 ($\downarrow$ 9.3%) 157.36 ($\downarrow$ 15.3%)
Z-score 53.94 ($\uparrow$ 14.7%) 114.85 ($\downarrow$ 56.2%) 153.61 ($\downarrow$ 51.3%) 99.88 ($\downarrow$ 15.6%) 105.57 ($\downarrow$ 43.2%)
@@ 4.4 Impact of Viewpoint
One interesting experiment is to look at the interplay between the viewpoint distribution of the datasets and its impact on the MPJPE error. In figs. 6 and 7, we present the MPJPE scores against the viewpoint distribution over the training and testing sets simultaneously. One common trend across all these figures is that whenever the viewpoint distribution disagrees with the training set’s viewpoint distribution, we observe a spike in error. This shows how much critical viewpoint is in 3D human pose estimation. Table 5 formally measures the inverse relationship between training data frequency and joint error by computing the Spearman correlation between the two, measured across patches on the viewpoint sphere measuring 5°5  ° $\times$ 10°10°10  ° degrees in elevation and azimuth, respectively. The correlation is always negative with high statistical significance. Notice how the p-value is lower when the evaluation is cross-dataset, indicating the variance in viewpoint distribution severely impacts performance. Please refer to Section A.2 for more figures.
@@
Table 5: Spearman correlation (which contained at least 5 training images and at least 5 test images) between training viewpoint distributions and test error (5°5  ° elev × 10°10°10  ° azim bins). Num = number of bins; sigma is the p-value represented as the number of standard deviations from random.
@@
[…]
380 SURREAL H36M -0.19 0.018 380.018380.018\,380.018 38 3.70
[…]
417 H36M 3DPW -0.20 0.003 7010.0037010.003\,7010.003 701 4.13
[…]
@@ 5 Conclusion and Future Work
We introduced a framework for evaluating cross-dataset performance of 2D-to-3D pose lifting networks. PoseBench3D enables standardized, reproducible evaluation and supports automatic benchmarking across multiple datasets. We apply PoseBench3D to evaluate 18 configurations of models, using checkpoints either obtained from prior work or retrained according to the original authors’ specifications. With more than 100 newly reported cross-dataset comparisons, we analyzed the results through several lenses and identify key factors—such as viewpoint distribution and normalization strategies—that significantly impact generalization in 3D human pose estimation.
@@
Appendix A Experiments
A.1 Example User-Supplied Config File
Listing 1: User-Supplied YAML Config File
# Unified YAML configuration
# General settings
dataset: [h36m, surreal, 3dpw, gpa]
keypoints: ["gt": ground truth]
checkpoint: "checkpoint"
evaluate: "/path_to_checkpoint.pt"
model_type: [JIT-Traced PyTorch File, ONNX TensorFlow File]
# Dataset locations
path_to_h36m: path_to_h36m
path_to_gpa: path_to_gpa
path_to_surreal_train: path_to_surreal_train
path_to_surreal_test: path_to_surreal_test
path_to_3dpw: path_to_3dpw
data_dir: path_to_data_directory
# Runtime & miscellaneous options
num_workers: 2-4
print_sample: [True, False]
per_joint_error: [True, False]
save_predictions: [True, False]
# Input tensor shape
input_shape:
batch_size: [1…N]
num_frames: 1
num_joints: [14, 16]
coordinates: 2
# Output tensor shape
output_shape:
batch_size: [1…N]
num_frames: 1
num_joints: [14, 16]
coordinates: 3
# Model-specific flags
model_info:
flattened_coordinate_model: [True, False]
output_3d: [Meters, Millimeters]
root_center_2d_test_input: [True, False]
normalize_2d_to_minus_one_to_one: [True, False]
trained_on_normalized_data: [True, False]
video_model: [True, False]
A.2 Cross-Dataset Evaluation Visualization
Figure 8: Protocol 1 visual results from table 2 shown in increasing average dataset error.
Figure 9: Protocol 2 visual results from table 3 shown in increasing average dataset error.
Figure 10: Per-joint errors for Protocols 1 and 2, averaged across all models on H36M.
Figure 11: Per-joint errors for Protocols 1 and 2, averaged across all models on GPA.
Figure 12: Per-joint errors for Protocols 1 and 2, averaged across all models on 3DPW.
Figure 13: Per-joint errors for Protocols 1 and 2, averaged across all models on SURREAL.
(a) H36M
(b) GPA
(c) 3DPW
(d) SURREAL
Figure 14: MPJPE score (SEM-GCN [42] trained on GPA) compared against viewpoint distribution in azimuth relative to the subject.
(a) H36M
(b) GPA
(c) 3DPW
(d) SURREAL
Figure 15: MPJPE score (SEM-GCN [42] trained on GPA) compared against viewpoint distribution in elevation relative to the subject.
(a) H36M
(b) GPA
(c) 3DPW
(d) SURREAL
Figure 16: MPJPE score (SEM-GCN [42] trained on 3DPW) compared against viewpoint distribution in azimuth relative to the subject.
(a) H36M
(b) GPA
(c) 3DPW
(d) SURREAL
Figure 17: MPJPE score (SEM-GCN [42] trained on 3DPW) compared against viewpoint distribution in elevation relative to the subject.
(a) H36M
(b) GPA
(c) 3DPW
(d) SURREAL
Figure 18: MPJPE score (SEM-GCN [42] trained on GPA) compared against viewpoint distribution in elevation relative to the subject.
(a) H36M
(b) GPA
(c) 3DPW
(d) SURREAL
Figure 19: MPJPE score (SEM-GCN [42] trained on SURREAL) compared against viewpoint distribution in azimuth relative to the subject.
(a) H36M
(b) GPA
(c) 3DPW
(d) SURREAL
Figure 20: MPJPE score (SEM-GCN [42] trained on SURREAL) compared against viewpoint distribution in elevation relative to the subject.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: Both the abstract and introduction accurately reflect the paper’s scope. We discuss the current friction and research gap, how we set out to address these issues, introduce a benchmarking evaluation environment to partner our approach, and provide results according to our proposed approach.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We discuss the limitation and future work in the conclusion section (section 5).
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: The paper does not contain theoretical results; instead, it focuses on the practical applications of human pose estimation.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: The paper outlines in detail the requirements necessary to reproduce our results. We mention all three components: an extensive overview of required datasets section 3.1, the models and architectures supported by the framework section 3.2, and the user-supplied configuration file for customization section A.1. We make our codebase available to the public for extension and reproduction of our work.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: Our framework is fully open-access, with a link to its repository provided. The repository includes detailed instructions for dataset setup, model import, and configuration file requirements needed to reproduce and extend our results.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We provide details about our experimental settings/details in the github repository linked with the paper. While we mention the specifics of our testing environment, our framework does not support training. For model training details and settings, please refer to the original work by the authors.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: We present the statistical significance of the evaluation results against viewpoint distribution through Spearman correlation (c.f. table 5).
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We present the compute resources in section 4.1.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: Our paper clearly details the code and resources used in this project to ensure both the reproducibility of our results and ease of extension for future work. All raw datasets are publicly available at the original sources provided by their respective authors. Additionally, we provide the post-processed data used in our experiments.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: We highlight the positive impact of advancing 3D Human Pose Estimation, particularly in rapidly growing fields as noted in the Introduction. We do not believe there are any significant negative societal impacts associated with improving 3D Human Pose Estimation.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: We expect that 3D Human Pose Estimation benchmarking and the datasets have a low risk for misuse.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We use four datasets, all of which are free for research purposes and credited with appropriate citations. The eighteen model configurations that we use for experiments or re-evaluation are licensed under one of the following: CC-BY-4.0, MIT, Apache License 2.0. Some of the models do not mention any citation. Nevertheless, we cited all the models used in our experiments.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We provide the original code repository, checkpoints, and data preparation script with proper documentation.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: Our work does not make use of crowdsourcing or human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: Our work does not make use of human subjects
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: Our work does not make use of LLMs in a way that is important to the core methods of this paper.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.

### v2

@@ 4.3 Impact of Z-Score Standardization
Z-score standardization is a common operation to normalize the data by zero-centering with the mean and scaling by the standard deviation. In the context of pose estimation, where datasets may originate from diverse distributions, such normalization indeed improves cross-dataset performance when the mean and standard deviation are known—though they are typically unavailable during inference time in real-world scenarios. Table˜4 shows the models [22] and [43] trained with H36M and normalized with the mean and standard deviation of the same dataset. For the other datasets, we used their respective mean and standard deviation to normalize them. We also include the unoptimized ($\diamond$) and PoseAug-optimized ($\dagger$) variants from Tab.˜2 for comparison, and also note the percentage improvement relative to the unoptimized variant in both cases. From all the cases, note that Z-score normalization performed with the test set’s mean and standard deviation improves cross-dataset generalization significantly for both models (e.g., SemGCN sees an improvement of more than 50% in 3DPW). PoseAug’s optimization adds novel poses to the training and secures a place right in the middle of the unoptimized variant and the Z-score standardized variant. This shows that models overfit on the distribution they are trained on—which can often be countered by increasing variation.
Table 4: MPJPE comparison of Martinez and SEM-GCN models across datasets. Percentage improvements relative to the baseline ($\diamond$) are in parentheses. Arrows in green indicate improvement over the Unoptimized (baseline), and red arrows indicate degradation from baseline.
@@
Unoptimized $\diamond$ 41.42 205.62 226.20 110.01 145.81
Optimized $\dagger$ 39.11 ($\downarrow$ 5.6%) 169.79 ($\downarrow$ 17.4%) 134.12 ($\downarrow$ 40.7%) 98.99 ($\downarrow$ 10.0%) 110.50 ($\downarrow$ 24.2%)
Z-score Normalization 52.37 ($\uparrow$ 26.4%) 104.39 ($\downarrow$ 49.3%) 141.10 ($\downarrow$ 37.6%) 81.64 ($\downarrow$ 25.8%) 94.88 ($\downarrow$ 34.9%)
@@
Unoptimized $\diamond$ 47.03 262.34 315.31 118.30 185.75
Optimized $\dagger$ 41.90 ($\downarrow$ 10.9%) 241.21 ($\downarrow$ 8.1%) 239.07 ($\downarrow$ 24.2%) 107.26 ($\downarrow$ 9.3%) 157.36 ($\downarrow$ 15.3%)
Z-score Normalization 53.94 ($\uparrow$ 14.7%) 114.85 ($\downarrow$ 56.2%) 153.61 ($\downarrow$ 51.3%) 99.88 ($\downarrow$ 15.6%) 105.57 ($\downarrow$ 43.2%)
@@ 4.4 Impact of Viewpoint
One interesting experiment is to look at the interplay between the viewpoint distribution of the datasets and its impact on the MPJPE error. In Figs.˜6 and 7, we present the MPJPE scores against the viewpoint distribution over the training and testing sets simultaneously. One common trend across all these figures is that whenever the viewpoint distribution disagrees with the training set’s viewpoint distribution, we observe a spike in error. This shows how much critical viewpoint is in 3D human pose estimation. Table˜5 formally measures the inverse relationship between training data frequency and joint error by computing the Spearman correlation between the two, measured across patches on the viewpoint sphere measuring 5° $\times$ 10° degrees in elevation and azimuth, respectively. The correlation is always negative with high statistical significance. Notice how the p-value is lower when the evaluation is cross-dataset, indicating the variance in viewpoint distribution severely impacts performance. Please refer to Figure˜14 and Figure˜15 for greater comparison of error rates across models and datasets.
@@
Table 5: Spearman correlation (which contained at least 5 training images and at least 5 test images) between training viewpoint distributions and test error (5° elev × 10° azim bins). Num = number of bins; sigma is the p-value represented as the number of standard deviations from random.
@@
377 3DPW GPA -0.45 1.1e181.110-18 9.78
380 3DPW H36M -0.30 2.4e072.410-07 6.10
388 3DPW SURREAL -0.64 1.1e471.110-47 16.38
304 3DPW 3DPW -0.44 2.8e142.810-14 8.55
377 SURREAL GPA -0.33 2.6e092.610-09 6.85
380 SURREAL H36M -0.19 0.018 380.018\,38 3.70
388 SURREAL SURREAL -0.47 1.2e211.210-21 10.55
304 SURREAL 3DPW -0.35 1.5e081.510-08 6.58
641 H36M GPA -0.36 3.4e193.410-19 9.78
939 H36M H36M -0.55 1.9e781.910-78 20.36
891 H36M SURREAL -0.61 7.5e967.510-96 22.90
417 H36M 3DPW -0.20 0.003 7010.003\,701 4.13
621 GPA GPA -0.48 1.3e361.310-36 13.71
738 GPA H36M -0.63 3.0e873.010-87 22.02
751 GPA SURREAL -0.68 6.8e1126.810-112 25.50
416 GPA 3DPW -0.39 1.3e141.310-14 8.58
4.5 Correlation between Viewpoint and Error
In Figs.˜6 and 7, we presented the error distribution curve with respect to the elevation and azimuth distribution separately. However, azimuth and elevation alone do not dictate the error—the number of training figures at a given point on the sphere dictates the error (i.e., both elevation and azimuth together). Therefore, we include additional contour plots showing the viewpoint and error distribution. Table˜5 showed a strong inverse correlation between the error and viewpoint distribution with high statistical significance. This is not observed in individual azimuth plots (e.g., Fig.˜6). In the contour plots (e.g., Figs.˜9(b), 11(b) and 13(b)), the viewpoint and error distribution are marked with green and red heatmaps. The X and Y axes represent azimuth and elevation, respectively. Whenever there is a trough in the viewpoint distribution—i.e. the training samples are under-sampled—the error goes up. This shows how much critical viewpoint is in 3D human pose estimation.
Figure 9: (a) Viewpoint distribution vs. MPJPE ($m$) error. (b) Contour plot showing the error (in red) and viewpoint distribution (in green) with the x and y axes as azimuth and elevation, respectively. For both figures, errors are obtained on the GPA test set from [43] trained with the H36M dataset.
Figure 11: (a) Viewpoint distribution vs. MPJPE ($m$) error. (b) Contour plot showing the error (in red) and viewpoint distribution (in green) with the x and y axes as azimuth and elevation, respectively. For both figures, errors are obtained on the 3DPW test set from [43] trained with the H36M dataset.
Figure 13: (a) Viewpoint distribution vs. MPJPE ($m$) error. (b) Contour plot showing the error (in red) and viewpoint distribution (in green) with the x and y axes as azimuth and elevation, respectively. For both figures, errors are obtained on the SURREAL test set from [43] trained with the H36M dataset.
Figure 14: Protocol 1 visual results from Tab.˜2 shown in increasing average dataset error.
Figure 15: Protocol 2 visual results from Tab.˜3 shown in increasing average dataset error.
@@ 5 Conclusion and Future Work
We have introduced a framework for evaluating cross-dataset performance of 2D-to-3D pose lifting networks. PoseBench3D enables standardized, reproducible evaluation and supports automatic benchmarking across multiple datasets. We apply PoseBench3D to evaluate 18 configurations of models, using checkpoints either obtained from prior work or retrained according to the original authors’ specifications. With more than 100 newly reported cross-dataset comparisons, we analyzed the results through several lenses and identify key factors—such as viewpoint distribution and normalization strategies—that significantly impact generalization in 3D human pose estimation.